OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? 54 27 81 62

OpenStudy (anonymous):

b + g = 89 b = 2g + 8 b = 89 - g b = 2g + 8 89 - g = 2g + 8 3g = 81 g = 27 b = 62

OpenStudy (anonymous):

The first line is the original 2 equations for the problem. The second line sets both equation into the form "b =" for upcoming substitution. Line 3: the substitution Line 4: simplification Line 5 more simplification Line 6 is determined by substitution back into equation #1.

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

ys thank you again

OpenStudy (anonymous):

uw